OpenStudy (anonymous):

determine the equation of the axis of symmetry. 8x^2+7x+6

OpenStudy (anonymous):

same as before this time \(-\frac{b}{2a}=-\frac{7}{2\times 8}=-\frac{7}{16}\)

OpenStudy (anonymous):

thank youuu againn